# Flow-based Conformal Prediction for Multi-dimensional Time Series

### v2

@@ 4.1 Marginal Coverage
We first establish that prediction sets generated by our method achieve exact non-asymptotic marginal coverage. This result follows from a fundamental property of flow: probability mass preservation under push-forward operations. When any measurable set is transformed through the push-forward operation of a flow, its probability mass is preserved. Lemma 4.3 formalizes this property and suffices to prove the exact non-asymptotic marginal coverage stated in Proposition 4.4.
@@ 4.2 Conditional Coverage
The true non-conformity scores $\{e_{i}\}_{i=1}^{T+1}$ are i.i.d.
@@
Lemma A.8 shows that bi-Lipschitz guided vector field results in bi-Lipschitz guided flow. Therefore, the vector field $u_{t}(x\mid h)$ can be modeled using neural network architectures that satisfy this assumption. For example, one can use invertible Residual Networks (iResNet) [behrmann2019invertible, chen2019residual] with smooth activation functions.
Assumption 4.8 (Lipschitz continuous of the CDF of the true non-conformity scores).
Assume that $F_{e}(u)$ is Lipschitz continuous with Lipschitz constant $L_{T+1}>0$, and that $F_{e}$ is strictly increasing in $u$.
@@
Lemma 4.10 (Convergence of empirical CDF of i.i.d. $\{e_{i}\}_{i=1}^{T}$).
Lemma 4.11 (Norm concentration of isotropic Gaussian random vectors).
$\max_{1\leq i\leq T}\|X_{i}\|\leq M_{T},$ (24)
Lemma 4.12 (Distance between the empirical CDF of $\{e_{i}\}_{i=1}^{T}$ and $\{\hat{e}_{i}\}_{i=1}^{T}$).
where $C=\sqrt{M_{T}L_{\psi^{-1}}\delta_{T}+L_{\psi^{-1}}^{2}\delta_{T}^{2}}$.
As a result of Lemma 23 and 4.12, Theorem 26 establishes the finite-sample bound for conditional coverage under i.i.d. non-conformity scores.
@@
Under Assumption 4.5, 4.6, 4.8, and 4.9, suppose the guided flow provides a sufficiently accurate approximation of the target distribution from the source distribution. With probability $1-\delta$, we have:
@@
Definition 4.14.
Definition 4.15.
@@
Assume that the sequence $\{e_{i}\}_{i=1}^{T+1}$ is strictly stationary and strongly mixing, with mixing coefficients satisfying $0<\sum_{k>0}\alpha(k)<M<\infty$.
Lemma 4.17 (Convergence of empirical CDF of stationary and strongly mixing $\{e_{i}\}_{i=1}^{T}$).
@@
4.6, 4.8, 4.9, and 4.16, suppose the guided flow provides a sufficiently accurate approximation of the target distribution from the source distribution. With probability $1-\delta$, we have:
@@
The bounds in Theorem 26 and Corollary 29 depend on the sample size $T$ and the estimation error $\delta_{T}$. Both bounds converge to $1-\alpha$ as $T\to\infty$, provided that $\delta_{T}=\mathcal{O}(T^{-a})$ for some $a>0$. Intuitively, with sufficiently large training data and an accurate base predictor $\hat{f}$, the conditional coverage is guaranteed. The condition on $\delta_{T}$ can be satisfied by a broad class of estimators. For example, sieve estimators based on general neural networks achieve $\delta_{T}=o_{p}(T^{-1/4})$ when $f$ is sufficiently smooth [chen1999improved]. The Lasso estimator and Dantzig selector achieve $\delta_{T}=o_{p}(T^{-1/2})$ when $f$ is a sparse high-dimensional linear model [bickel2009simultaneous].
@@ 5 Experiments
Table 1: Average empirical coverage and prediction sets sizes obtained by FCP and all baselines on three real-world datasets, evaluated under different base predictors and varying outcome dimensions $d_{y}$. Reported values represent the average and standard deviation over five independent experiments. The target confidence level was set to 0.95. Results with average empirical coverage below the target confidence level are grayed out, and the smallest prediction set sizes, excluding the grayed-out results, are highlighted in bold.
@@
Wind LOO Bootstrap FCP $0.951_{\pm.018}$ $\textbf{0.88}_{\pm.089}$ $0.953_{\pm.006}$ $3.43_{\pm 1.37}$ $0.956_{\pm.010}$ $19.4_{\pm 10.2}$
@@
LSTM FCP $0.952_{\pm.054}$ $\textbf{1.18}_{\pm.215}$ $0.957_{\pm.022}$ $10.8_{\pm 1.05}$ $0.953_{\pm.056}$ $\textbf{2.48}\times\textbf{10}^{\textbf{3}}_{\pm 669}$
@@
Local Ellipsoid $0.978_{\pm.043}$ $10.5_{\pm 6.97}$ $1.000_{\pm.000}$ $354.4_{\pm 406.8}$ $1.000_{\pm.000}$ $2.63\times 10^{5}_{\pm 2.70\times 10^{5}}$
Empirical Copula $0.983_{\pm.035}$ $14.2_{\pm 8.19}$ $1.000_{\pm.000}$ $494.5_{\pm 196.1}$ $1.000_{\pm.000}$ $4.46\times 10^{5}_{\pm 9.82\times 10^{4}}$
Gaussian Copula $0.983_{\pm.035}$ $14.1_{\pm 8.18}$ $1.000_{\pm.000}$ $499.1_{\pm 189.5}$ $1.000_{\pm.000}$ $5.24\times 10^{5}_{\pm 1.89\times 10^{5}}$
@@
TFT $0.407_{\pm.065}$ $0.292_{\pm.089}$ $189_{\pm.306} 0.0$$70_{\pm.031} 0.0$$90_{\pm.007} 0.0$$009_{\pm.007}$
DeepAR $0.443_{\pm.095}$ $0.308_{\pm.088}$ $197_{\pm.054} 0.0$$70_{\pm.030} 0.0$$90_{\pm.028} 0.0$$004_{\pm.003}$
LSTM FCP $0.968_{\pm.022}$ $0.859_{\pm.075}$ $0.966_{\pm.022}$ $\textbf{1.05}_{\pm.111}$ $0.950_{\pm.010}$ $\textbf{1.82}_{\pm.287}$
@@
DeepAR $0.386_{\pm.065}$ $0.266_{\pm.069}$ $211_{\pm.056} 0.0$$6_{\pm.017} 0.0$$90_{\pm.009} 0.0$$003_{\pm.001}$
Solar LOO Bootstrap FCP $0.957_{\pm.007}$ $1.48_{\pm.292}$ $0.969_{\pm.003}$ $4.18_{\pm.597}$ - -
@@
For notational convenience, we refer to our method as FCP, which stands for Flow-based Conformal Prediction. We use MLP with Softplus activation to model the guided vector field and concatenate the guidance and time with the input for the MLP. dopri5 [dormand1980family] at absolute and relative tolerances of 1e-5 is used to solve all ODEs. A grid search is conducted to select the optimal hyperparameters for FCP. To determine an appropriate sample size $N$, we compute the relative standard error (SE) of the Jacobian determinants of $\psi$, defined as $\textrm{SE}\left(\{\det J_{\psi}(x_{j}\mid h)\}_{j=1}^{N}\right)/\textrm{Avg}\left(\{\det J_{\psi}(x_{j}\mid h)\}_{j=1}^{N}\right)$, then choose the smallest $N$ such that the average relative SE across all $h$ falls below 0.01. The source code for FCP is available at github.com/Jayaos/flow_cp.
Baselines.
We evaluate FCP against several conformal prediction methods covering various existing approaches: MultiDimSPCI [xu2024conformal], OT-CP [thurinoptimal], CONTRA [fang2025contra], conformal prediction using local ellipsoids [messoudi2022ellipsoidal], CopulaCPTS [sun2022copula], and conformal prediction using empirical and Gaussian copulas [messoudi2021copula]. We also include two widely used probabilistic time series forecasting methods as baselines: Temporal Fusion Transformer (TFT) [lim2021temporal] and DeepAR [salinas2020deepar]. Although TFT and DeepAR are originally developed for time series with univariate outcomes, we adapt them to our multi-dimensional setting by constructing independent copulas using the predicted intervals for each output dimension. Additional details and setup of the baselines are provided in Appendix B.
Datasets and base predictor.
We evaluated FCP and baselines on three real-world time series datasets: wind, traffic, and solar datasets. For the wind and traffic datasets, we randomly selected $d_{y}\in\{2,4,8\}$ locations to construct five sequences of $d_{y}$-dimensional time series. For the solar dataset, we use $d_{y}\in\{2,4\}$ and similarly construct five sequences. Additional dataset details are provided in Appendix C. Base predictor $\hat{f}$ is required to provide a point prediction $\hat{y}$. We used two types of base predictors for each dataset: (1) leave-one-out (LOO) bootstrap ensemble of 15 multivariate linear regressors, and (2) recurrent neural network (RNN) with long short-term memory (LSTM) units [hochreiter1997long]. Since the RNN base predictor requires part of the sequence for training, whereas the LOO bootstrap predictor can leverage the full sequence, the effective sequence length available for evaluation varies by predictor. Each base predictor was trained independently for every sequence. For the RNN base predictor, the first 50% of each sequence was allocated for training, and predictions were made for the remaining 50%, which served as the evaluation sequence. Within this evaluation sequence, the first 80% was used as a training set, and the final 20% was evenly divided into validation and test sets. Since FCP does not require a calibration set to construct prediction sets, the validation set was used for model selection during training. To ensure fair evaluation in terms of data utilization, we combined the training and validation sets into a single calibration set for methods that require a calibration set. The specific data utilization scheme for each baseline is detailed Appendix B.
@@
where $\mathcal{D}_{\text{test}}$ denotes the test set. The average prediction set size is computed by averaging the sizes of $\hat{C}_{i}$ over the test set, with the specific definition of the set size depending on the geometric form of each method.
@@ Results.
Table 1 presents the results of experiments on three real-world datasets. FCP consistently obtained smaller prediction sets than all baselines while maintaining the target coverage. The performance gains of FCP were especially notable for higher outcome dimensions, showing significantly smaller prediction set sizes with lower variability. Moreover, FCP maintained stable coverage across varying $d_{y}$, whereas baseline methods often suffered from undercoverage or from overcoverage coupled with either overly contracted or excessively inflated prediction sets. In particular, methods relying on the exchangeability assumption often exhibited severe coverage errors and highly unstable prediction set sizes.
MultiDimSPCI and CP using local ellipsoids generally showed good performance. In particular, on the solar dataset, CP with local ellipsoids achieved performance comparable to FCP. This is possibly due to their ability to capture temporal or local correlations, respectively. OT-CP and CONTRA also performed well in certain experiments, indicating some potential to adapt beyond the exchangeability assumption. We observed that increasing the guidance scale $w$ often reduced the prediction set size, though at the cost of slightly lower coverage. In practice, an effective range for $w$ was typically between 1 and 1.5 across our experiments.
Table 2: Average empirical coverage and prediction set sizes obtained by FCP and FCP without the encoder on the wind dataset, evaluated under different base predictors and varying outcome dimensions $d_{y}$. The target confidence level was set to 0.95.
@@
LOO Bootstrap FCP (MLP) $0.951_{\pm.018}$ $0.88_{\pm.089}$ $0.953_{\pm.006}$ $3.43_{\pm 1.37}$ $0.956_{\pm.010}$ $19.4_{\pm 10.2}$
FCP (MLP w/o Encoder) $0.948_{\pm.023}$ $1.13_{\pm.193}$ $0.964_{\pm.005}$ $3.99_{\pm 1.03}$ $0.964_{\pm.010}$ $35.3_{\pm 14.0}$
LSTM FCP (MLP) $0.952_{\pm.054}$ $1.18_{\pm.215}$ $0.957_{\pm.022}$ $10.8_{\pm 1.05}$ $0.953_{\pm.056}$ $2.48\times 10^{3}_{\pm 669}$
FCP (MLP w/o Encoder) $0.965_{\pm.011}$ $1.92_{\pm.367}$ $0.957_{\pm.014}$ $12.2_{\pm 15.0}$ $0.935_{\pm.007}$ $5.55\times 10^{3}_{\pm 7.47\times 10^{3}}$
Ablation study.
We conduct an ablation study to assess the impact of the encoder. Specifically, we evaluate FCP with and without the encoder, where in the latter case the guidance $h$ is replaced by the concatenation of the feature at time $i$ and residual at time i1i-1. Table 2 reports the average empirical coverage and prediction set sizes of FCP with and without the encoder on the wind dataset. We observe that removing the encoder led to less stable coverage and noticeably larger prediction set sizes.
Since the conditional coverage bound of FCP relies on the bi-Lipschitz flow assumption (Assumption 4.6), we conduct an additional experiment using iResNet [behrmann2019invertible] to model the vector field, ensuring this assumption is satisfied. Table 8 reports the average empirical coverage and prediction set sizes of FCP with MLP and iResNet across the three datasets with varying $d_{y}$. We observe that imposing bi-Lipschitzness in the vector field did not negatively affect either coverage or prediction set size.

### v3

@@ 4.1 Marginal Coverage
We first establish that prediction sets generated by our method achieve exact non-asymptotic marginal coverage. This result follows from a fundamental property of flow: probability mass preservation under push-forward operations. Lemma 4.3 formalizes this property and suffices to prove the exact non-asymptotic marginal coverage stated in Proposition 4.4.
@@ 4.2 Conditional Coverage
The true non-conformity scores $\{e_{i}\}_{i=1}^{T}$ are i.i.d.
@@
Lemma A.7 shows that bi-Lipschitz guided vector field results in bi-Lipschitz guided flow. Consequently, we can model the guided vector field to satisfy this assumption. For example, one may use invertible Residual Networks (iResNet) [4, 9] with smooth activation functions.
Assumption 4.8 (Lipschitz continuous of the CDF of true non-conformity scores).
$F_{e}(u)$ is Lipschitz continuous with Lipschitz constant $L_{T+1}>0$, and that $F_{e}$ is strictly increasing in $u$.
@@
As a result of Lemma 80 and A.14, Theorem 23 establishes the finite-sample bound for conditional coverage under i.i.d. non-conformity scores.
@@
Under Assumption 4.5, 4.6, 4.8, and 4.9, suppose the guided flow provides a sufficiently accurate approximation of the target distribution. With probability $1-\delta$, we have:
@@
Next, we present a finite-sample bound on conditional coverage under the assumption of stationary and strongly mixing non-conformity scores.
@@
$\{e_{i}\}_{i=1}^{T}$ are strictly stationary and strongly mixing, with mixing coefficients satisfying $0<\sum_{k>0}\alpha(k)<M<\infty$.
@@
Under Assumption 4.6, 4.8, 4.9, and 4.11, suppose the guided flow provides a sufficiently accurate approximation of the target distribution. With probability $1-\delta$, we have:
@@
The bounds in Theorem 23 and Corollary 24 depend on the sample size $T$ and the estimation error $\delta_{T}$. Both bounds converge to $1-\alpha$ as $T\to\infty$, provided that $\delta_{T}=\mathcal{O}(T^{-a})$ for some $a>0$. Intuitively, this implies the conditional coverage is guaranteed with sufficiently large training data and an accurate base predictor $\hat{f}$. The condition on $\delta_{T}$ can be satisfied by a broad class of estimators. For example, sieve estimators based on general neural networks achieve $\delta_{T}=o_{p}(T^{-1/4})$ when $f$ is sufficiently smooth [10]. The Lasso estimator and Dantzig selector achieve $\delta_{T}=o_{p}(T^{-1/2})$ when $f$ is a sparse high-dimensional linear model [5].
@@ 5 Experiments
For notational convenience, we refer to our method as FCP, which stands for Flow-based Conformal Prediction. The guided vector field is modeled by an MLP with Softplus activations. We concatenate the guidance and time with the input for the MLP. dopri5 [13] at absolute and relative tolerances of $10^{-5}$ is used to solve ODEs. A grid search is conducted to select the optimal hyperparameters for FCP. To determine an appropriate sample size $N$, we compute the relative standard error (SE) of the Jacobian determinants of $\psi$, defined as $\textrm{SE}\left(\{\det J_{\psi}(x_{j}\mid h)\}_{j=1}^{N}\right)/\textrm{Avg}\left(\{\det J_{\psi}(x_{j}\mid h)\}_{j=1}^{N}\right)$, then choose the smallest $N$ such that the average relative SE across all $h$ falls below 0.01. Additional details on hyperparameter selection and the choice of $N$ are provided in Appendix B.
Baselines.
We evaluate FCP against several conformal prediction methods: MultiDimSPCI [50], OT-CP [44], CONTRA [16], conformal prediction using local ellipsoids [31], CopulaCPTS [43], and conformal prediction using empirical and Gaussian copulas [30]. We also include two widely used probabilistic time series forecasting methods as baselines: Temporal Fusion Transformer (TFT) [28] and DeepAR [38]. Although TFT and DeepAR are originally developed for time series with univariate outcomes, we adapt them to our multi-dimensional setting by constructing independent copulas using the predicted intervals for each output dimension. Further details on the baseline methods and their experimental setup are provided in Appendix B.
Datasets and base predictor.
We evaluate FCP and baselines on three real-world time series datasets: wind, traffic, and solar datasets. Additional details on the datasets are provided in Appendix C. For the wind and traffic datasets, we randomly select $d_{y}\in\{2,4,8\}$ locations to construct five sequences of $d_{y}$-dimensional time series. For the solar dataset, we use $d_{y}\in\{2,4\}$ and similarly construct five sequences. Base predictor $\hat{f}$ is required to provide a point prediction $\hat{y}$. We use two types of base predictors: (1) leave-one-out (LOO) bootstrap ensemble of 15 multivariate linear regressors, and (2) recurrent neural network (RNN) with long short-term memory (LSTM) units [22]. Since the RNN base predictor requires part of the sequence for training, whereas the LOO bootstrap predictor can leverage the full sequence, the effective sequence length available for evaluation varies by base predictor. The base predictor is trained independently for each of the five constructed sequences. For the RNN base predictor, the first 50% of each sequence was allocated for training, and predictions were made for the remaining 50%, which served as the evaluation sequence. Within this evaluation sequence, the first 80% was used as a training set, and the final 20% was evenly divided into validation and test sets. Since FCP does not require a calibration set to construct prediction sets, the validation set was used for model selection during training. To ensure fair evaluation in terms of data utilization, we combined the training and validation sets into a single calibration set for methods that require a calibration set.
Table 1: Average empirical coverage and prediction sets size obtained by FCP and all baselines on three real-world datasets, evaluated under different base predictors and varying outcome dimensions $d_{y}$. Reported values represent the average and standard deviation over five independent experiments conducted on five constructed sequences. The target confidence level is set to 0.95. Results with average empirical coverage below the target confidence level are grayed out, and the smallest prediction set sizes, excluding the grayed-out results, are highlighted in bold.
@@
Wind LOO Bootstrap FCP $0.951_{\pm.018}$ $\textbf{0.88}_{\pm.089}$ $0.953_{\pm.006}$ $\textbf{3.43}_{\pm 1.37}$ $0.956_{\pm.010}$ $\textbf{19.4}_{\pm 10.2}$
@@
LSTM FCP $0.952_{\pm.054}$ $\textbf{1.18}_{\pm.215}$ $0.957_{\pm.022}$ $\textbf{10.8}_{\pm 1.05}$ $0.953_{\pm.056}$ $\textbf{2.48}\times\textbf{10}^{\textbf{3}}_{\pm 669}$
@@
Local Ellipsoid $0.978_{\pm.043}$ $10.5_{\pm 6.97}$ $1.0_{\pm.000}$ $354.4_{\pm 406.8}$ $1.0_{\pm.000}$ $2.63\times 10^{5}_{\pm 2.70\times 10^{5}}$
Empirical Copula $0.983_{\pm.035}$ $14.2_{\pm 8.19}$ $1.0_{\pm.000}$ $494.5_{\pm 196.1}$ $1.0_{\pm.000}$ $4.46\times 10^{5}_{\pm 9.82\times 10^{4}}$
Gaussian Copula $0.983_{\pm.035}$ $14.1_{\pm 8.18}$ $1.0_{\pm.000}$ $499.1_{\pm 189.5}$ $1.0_{\pm.000}$ $5.24\times 10^{5}_{\pm 1.89\times 10^{5}}$
@@
TFT $0.407_{\pm.065}$ $0.292_{\pm.089}$ $189_{\pm.306} 0.0$$7_{\pm.031} 0.0$$9_{\pm.007} 0.0$$009_{\pm.007}$
DeepAR $0.443_{\pm.095}$ $0.308_{\pm.088}$ $197_{\pm.054} 0.0$$7_{\pm.030} 0.0$$9_{\pm.028} 0.0$$004_{\pm.003}$
LSTM FCP $0.968_{\pm.022}$ $\textbf{0.859}_{\pm.075}$ $0.966_{\pm.022}$ $\textbf{1.05}_{\pm.111}$ $0.950_{\pm.010}$ $\textbf{1.82}_{\pm.287}$
@@
DeepAR $0.386_{\pm.065}$ $0.266_{\pm.069}$ $211_{\pm.056} 0.0$$6_{\pm.017} 0.0$$9_{\pm.009} 0.0$$003_{\pm.001}$
Solar LOO Bootstrap FCP $0.957_{\pm.007}$ $\textbf{1.48}_{\pm.292}$ $0.969_{\pm.003}$ $\textbf{4.18}_{\pm.597}$ - -
@@
where $\mathcal{D}_{\text{test}}$ denotes the test set. The average prediction set size is computed by averaging the sizes of $\hat{C}_{i}$ over the test set, with the specific definition of the set size depending on each method.
@@ Results.
Table 1 presents the results of experiments on three real-world datasets. FCP consistently obtains smaller prediction sets than all baselines while maintaining the target coverage. The performance gains of FCP are especially notable for higher outcome dimensions, showing significantly smaller prediction set sizes with lower variability. Moreover, FCP maintains stable coverage across varying $d_{y}$, whereas baseline methods often suffer from undercoverage or overcoverage coupled with either overly contracted or inflated prediction sets. In particular, methods relying on the exchangeability assumption often exhibit severe miscoverages and unstable prediction set sizes.
MultiDimSPCI and CP using local ellipsoids generally show good performance. In particular, on the solar dataset, CP with local ellipsoids achieves performance comparable to FCP. This is possibly due to their ability to capture temporal or local correlations, respectively. OT-CP and CONTRA also perform well in certain experiments, indicating some potential to adapt beyond the exchangeability assumption. We observe that increasing the guidance scale $w$ often reduces the prediction set size, though at the cost of slightly lower coverage. In practice, an effective range for $w$ is typically between 1 and 1.5 across our experiments.
Ablation study.
We conduct an ablation study to assess the impact of the encoder. Specifically, we evaluate FCP with and without the encoder, where in the latter case the guidance is replaced by the concatenation of the feature at time $i$ and residual at time i1i-1. Table 2 reports the average empirical coverage and prediction set sizes of FCP with and without the encoder on the wind dataset. We observe that removing the encoder led to less stable coverage and noticeably larger prediction set sizes.
Since the conditional coverage bound of FCP relies on the bi-Lipschitz flow assumption (Assumption 4.6), we conduct an additional ablation study using iResNet [4] to model the guided vector field, ensuring this assumption is satisfied. Table 7 reports the average empirical coverage and prediction set sizes of FCP with MLP and iResNet across the three datasets with varying $d_{y}$. We observe that imposing bi-Lipschitzness in the vector field does not degrade coverage or prediction set size.
Table 2: Average empirical coverage and prediction sets size obtained by FCP with and without the encoder on the wind dataset, evaluated under different base predictors and varying outcome dimensions $d_{y}$. Reported values represent the average and standard deviation over five independent experiments conducted on five constructed sequences. The target confidence level is set to 0.95.
@@
LOO Bootstrap FCP with Encoder $0.951_{\pm.018}$ $0.88_{\pm.089}$ $0.953_{\pm.006}$ $3.43_{\pm 1.37}$ $0.956_{\pm.010}$ $19.4_{\pm 10.2}$
FCP w/o Encoder $0.948_{\pm.023}$ $1.13_{\pm.193}$ $0.964_{\pm.005}$ $3.99_{\pm 1.03}$ $0.964_{\pm.010}$ $35.3_{\pm 14.0}$
LSTM FCP with Encoder $0.952_{\pm.054}$ $1.18_{\pm.215}$ $0.957_{\pm.022}$ $10.8_{\pm 1.05}$ $0.953_{\pm.056}$ $2.48\times 10^{3}_{\pm 669}$
FCP w/o Encoder $0.965_{\pm.011}$ $1.92_{\pm.367}$ $0.957_{\pm.014}$ $12.2_{\pm 15.0}$ $0.935_{\pm.007}$ $5.55\times 10^{3}_{\pm 7.47\times 10^{3}}$